OpenStudy (anonymous):

if x/2= 3/x+1- 1/2 then x= ?

OpenStudy (anonymous):

x=-2 x=3

OpenStudy (saifoo.khan):

@MMH123 , do you get how he did it?

OpenStudy (anonymous):

What was your answer?

OpenStudy (anonymous):

no

OpenStudy (saifoo.khan):

\[\frac x2= \frac{3}{x+1}- \frac12\]is it that?

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

Okay, let's solve the right side first, \[\frac x2 = \frac{3 -1(x+1)}{2(x+1)}\]\[\frac x2 = \frac{3 - x - 1}{2x+2}\]

OpenStudy (saifoo.khan):

\[\frac x2 = \frac{2 - x}{2x + 2}\]Cross multiply.. \[x(2x+2) = 2(2-x)\]Can you solve now?

OpenStudy (anonymous):

yeah thanks

OpenStudy (saifoo.khan):

Welcome.

OpenStudy (asnaseer):

@saifoo.khan you made a slight error up there - the "3 -" should be "6 -"...

OpenStudy (saifoo.khan):

OH! :/ Thanks @asnaseer Bhai.

OpenStudy (asnaseer):

yw

OpenStudy (saifoo.khan):

@MMH123 , there's a mistake..

OpenStudy (anonymous):

what's the mistake?

OpenStudy (saifoo.khan):

let me do it again..

OpenStudy (saifoo.khan):

Do you mind reposting it?

OpenStudy (saifoo.khan):

nvm.. im solving here.

OpenStudy (saifoo.khan):

\[\frac x2= \frac {3}{x+1}- \frac12\] Take the LCM 2(x+1) \[\frac x2 = \frac{3(2) - 1(x+1)}{2(x+1)}\]can u solve from here?

OpenStudy (anonymous):

yeah

OpenStudy (saifoo.khan):

Cool.

OpenStudy (anonymous):

i cant solve it from there

OpenStudy (saifoo.khan):

\[\frac x2 = \frac{6 - x - 1}{2x+2}\]Now?

OpenStudy (anonymous):

yeah thanks